OpenStudy (smileyxl3):

For the function defined by f(x)={x^2, x≤1} evaulate f(0)

OpenStudy (smileyxl3):

the function looks like this

OpenStudy (raden):

because 0 < 1 so chose f(x) = x^2 therefore the value of f(0) = 0^2 = 0

OpenStudy (smileyxl3):

@RadEn actually the function is the one I attached.

OpenStudy (raden):

yes, i used that

OpenStudy (smileyxl3):

How would you graph this?

OpenStudy (raden):

there are 2 choices which we have to choice one of them, depends for interval x is given. look at each interval for x's in defined of that function. if x < or equal 1, so select the 1st one. if x > 1 select the 2nd one. while the question asking f(0), the value of f(x) when x = 0. (notices that 0 < 1, right ?) because x < 1, so we select the 1st one as the function of f(x), it is f(x) = x^2 therefore f(0) = 0^2 = 0 hopes make sense